OpenStudy (tootsi123):

Identify the expression as an equation containing fractions, a fractional equation, or a proportion.

OpenStudy (tootsi123):

\[\frac{ x }{ 12 }=\frac{ 17 }{ 3 }-\frac{ 2x }{ 9 }-5+\frac{ 3x }{ 12 }\]

OpenStudy (tootsi123):

@sleepyjess @Luigi0210

OpenStudy (tootsi123):

and could this please be explained to me please?